OpenStudy (anonymous):

let f(x)=2/t...compute f(t)-f(a)/t-a...can any1 answer this step by step 4 me?

OpenStudy (anonymous):

lets compute f(t) and f(a) separately: \[f(t) = \frac{2}{t}, f(a) = \frac{2}{a}\] Now lets use that formula given in the problem: \[\frac{f(t)-f(a)}{t-a} = \frac{\frac{2}{t}-\frac{2}{a}}{t-a} = \frac{\frac{2a-2t}{at}}{t-a} = \frac{2(a-t)}{at(t-a)} = \frac{-2(t-a)}{at(t-a)} = \frac{-2}{at}\]

OpenStudy (anonymous):

how do u get at(t-a) on the bottom?

OpenStudy (anonymous):

the bottom (t-a) comes straight from the formula in the question

OpenStudy (anonymous):

i guess my real question is how does at come 2 b in front of (t-a)?

OpenStudy (anonymous):

oh oh the 'at' part. my bad. let me show a few more steps in that part: \[\frac{\frac{2a-2t}{at}}{t-a} = \frac{\frac{2a-2t}{at}}{t-a}*\frac{at}{at} = \frac{\frac{2a-2t}{at}*at}{at(t-a)} = \frac{2a-2t}{at(t-a)}\]

OpenStudy (anonymous):

you multiply the top and bottom of the fraction by 'at' to get rid of that at in the bottom of the top fraction.

OpenStudy (anonymous):

so in any situation where u have a fraction on top like this and u want to get rid of it u multi. the denom. to the top n bottom?

OpenStudy (anonymous):

yes, thats correct :)

OpenStudy (anonymous):

aha,gotcha.thanx dude!

myininaya (myininaya):

remember \[\frac{\frac{f}{g}}{\frac{h}{k}}=\frac{f}{g} \div \frac{h}{k}=\frac{f}{g} \times \frac{k}{h}\]